# From Timestamps to Versions: Version AoI in Single- and Multi-Hop Networks

### v1

@@ Abstract
Timely and informative data dissemination in communication networks is essential for enhancing system performance and energy efficiency, as it reduces the transmission of outdated or redundant data. Timeliness metrics, such as Age of Information (AoI), effectively quantify data freshness; however, these metrics fail to account for the intrinsic informativeness of the content itself. To address this limitation, content-based metrics have been proposed that combine both timeliness and informativeness. Nevertheless, existing studies have predominantly focused on evaluating average metric values, leaving the complete distribution—particularly in multi-hop network scenarios—largely unexplored. In this paper, we provide a comprehensive analysis of the stationary distribution of the Version Age of Information (VAoI), a content-based metric, under various scheduling policies, including randomized stationary, uniform, and threshold-based policies, with transmission constraints in single-hop and multi-hop networks. We derive closed-form expressions for the stationary distribution and average VAoI under these scheduling approaches. Furthermore, for threshold-based scheduling, we analytically determine the optimal threshold value that minimizes VAoI and derive the corresponding optimal VAoI in closed form. Numerical evaluations verify our analytical findings, providing valuable insights into leveraging VAoI in the design of efficient communication networks.
@@ I Introduction
Efficient data management is a critical requirement for ensuring optimal performance in communication networks across a wide range of applications, from single-hop IoT monitoring systems to multi-hop satellite-based networks. As the volume of data generated by these networks increases, transmitting all data indiscriminately, without considering its semantic significance or task-specific utility, becomes increasingly unsustainable. Such an approach results in excessive consumption of critical resources, including energy and bandwidth, ultimately compromising system practicality and degrading overall performance. To address these challenges, there is an urgent need for network management approaches that optimize data transmission by leveraging goal-oriented semantic metrics; that is, by delivering the most timely and informative data within a constrained frequency of data transmissions [1].
The AoI [2] is a widely used semantic metric that quantifies the freshness of information in status update systems as the time elapsed since the generation of the most recently received data. AoI-aware scheduling effectively minimizes staleness by adapting transmissions to source data arrivals and network service times. However, AoI captures freshness solely through data timestamps, without accounting for actual changes in the source content. As a result, simply refreshing timestamps may fail to deliver new information, and avoiding such redundant updates can reduce data transmission and energy consumption.
To address this limitation, content-based metrics such as Age of Incorrect Information (AoII) [3] and VAoI [4] have been introduced. AoII adds a distortion-aware dimension by measuring the staleness of incorrect information—specifically when the receiver’s content deviates from the source—unlike AoI, which treats both correct and incorrect data uniformly. However, AoII requires precise knowledge of the content or state of information at both source and destination for comparison, which is practical only when the state space is small and fully modeled, with all transitions known. In many real-world applications, such complete knowledge of source content and transitions may not be available. In such cases, VAoI offers a more practical, content-based metric by focusing solely on content changes at the source, where data evolve through successive, non-reverting versions. This requires minimal knowledge: at any time, either a new or the previous version exists, and the receiver must track these versions as timely as possible. Defined as the number of versions by which the receiver lags behind the source, VAoI further improves upon AoI by replacing timestamps with version numbers, thereby eliminating the challenging requirement of clock synchronization between the transmitter and receiver. It is computed simply by comparing the receiver’s stored version with the source’s current version.
@@ II Related Works
Several studies have investigated the distributions of AoI and Peak AoI (PAoI) in continuous-time systems using queueing theory [5, 6, 7, 8, 9, 10, 11, 12, 13, 14], while other studies have examined discrete-time settings [15, 16, 17, 18, 19]. Notably, [15] derives general expressions for the stationary distributions and generating functions of AoI and PAoI in discrete-time single-server queues under various disciplines, along with methods for evaluating nonlinear age functions. Extending stochastic hybrid system techniques to discrete time, [16] models AoI and packet age as a two-dimensional Markov process in bufferless queues with Bernoulli arrivals. A matrix-analytic framework based on quasi-birth–death processes is proposed in [17] to obtain exact per-source AoI and PAoI distributions in multi-source IoT systems with discrete phase-type service times under various queueing disciplines. The study in [18] investigates age-optimal packet scheduling with delayed feedback and long-term resource constraints, providing closed-form benchmarks for random and deterministic policies. Furthermore, [19] analyzes AoI and PAoI in multi-source Ber/Geo/1/1 systems under preemptive and non-preemptive policies, deriving closed-form expressions for both distributions and averages.
@@
Regarding multi-hop networks, [25] derives the distribution of discrete-time AoI in $N$-hop systems with time-invariant packet loss through recursive formulations, while [26] studies continuous-time AoI and PAoI distributions in two-hop scenarios. However, most existing research on multi-hop networks, including [27, 28, 29, 30, 31, 32, 33], primarily focuses on average metrics.
@@ III System Model
We consider a communication network for transmitting data from a source to a destination node, which may be separated by either a single hop or multiple hops. The data at the source are constantly sampled from an information source and then transmitted according to an update policy. The update policy schedules each data sample or update to be either transmitted or skipped, while satisfying a constraint on the long-term average transmission rate. Specifically, the average update rate must not exceed a predefined limit. We assume a slotted time axis $t\in\{0,1,2,\cdots\}$, and our objective is to investigate the stochastic distribution of the VAoI at the destination node in both single-hop and multi-hop scenarios under various rate-constrained policies. The details of the system model are explained below.
@@
Version Age of Information: We adopt the VAoI as the performance metric that captures both the timeliness and relevance of information. Unlike the AoI, which measures the time elapsed since the generation timestamp $u(t)$ of the freshest received update, $\Delta^{AoI}(t)\!=\!t\!-\!u(t),$ the VAoI quantifies how many versions the receiver lags behind the information source. At time slot $t$, the VAoI, denoted $\Delta(t)$, is defined as $\Delta(t)\!=\!V_{S}(t)\!-\!V_{R}(t)$, where $V_{S}(t)$ is the version index at the information source and $V_{R}(t)$ the version stored at the receiver. Fig. 4 shows the evolution of AoI and VAoI over discrete slots. A successful update occurs at $t\!=\!3$, causing AoI and VAoI to drop to 111 and 0, respectively, since the update is one slot old and no new versions have been generated. Between $t\!=\!3$ and the next update at $t\!=\!12$, the AoI grows linearly to 999, while the VAoI reaches 444, reflecting the generation of four new versions at the information source.
Fig. 5 illustrates VAoI dynamics in a two-hop network. The information source generates new versions at time slots 0, 222, 333, 666, and 888, with its version index $V_{S}(t)$ increasing by one at the start of each subsequent slot. Node 0 holds the most recent version available in the network; hence, its VAoI is always zero, i.e., $V_{0}(t)\!=\!V_{S}(t)$. The versions at node 111, denoted by $V_{1}(t)$, are updated according to the update policy and the transmissions from node 0. If a transmission at time $t$ succeeds, then $V_{1}(t\!+\!1)\!=\!V_{0}(t)$; otherwise, $V_{1}(t\!+\!1)\!=\!V_{1}(t)$. The corresponding VAoI, $\Delta_{1}(t)\!=\!V_{S}(t)\!-\!V_{1}(t)$, is listed in parentheses in the second row of the table and plotted in blue. Similarly, node 222 stores versions $V_{2}(t)$ received from node 111, which transmits in every slot. [^1] In practice, if feedback channels are available between relay nodes, retransmissions of already delivered versions can be avoided; otherwise, relay nodes transmit continuously each time slot. In both cases, however, the presented VAoI analysis remains valid. Here, $V_{2}(t\!+\!1)\!=\!V_{1}(t)$ if the transmission at $t$ succeeds; otherwise, $V_{2}(t\!+\!1)\!=\!V_{2}(t)$. The VAoI at node 222, $\Delta_{2}(t)\!=\!V_{S}(t)\!-\!V_{2}(t)$, is shown in parentheses in the third row of the table and plotted in green in Fig. 5.
@@ IV Analysis of VAoI in a Single-hop Setup
We analyze the VAoI in a single-hop status update system under the three aforementioned update policies using a dtmc model. The VAoI in the next time slot is stochastically determined by the current VAoI, the update policy, and system dynamics, including the version generation and channel success processes.
@@
where $\sum_{n=0}^{\infty}\mu_{n}=1$. Here, $P_{jn}$ denotes the transition probability from state $j$ to state $n$, i.e., $P_{jn}=\mathbb{P}\left(\Delta(t+1)=n\mid\Delta(t)=j\right)$. The steady-state distribution of a dtmc exists if it is irreducible, i.e., every state can be reached from every other state, and positive recurrent, i.e., the expected return time to each state is finite [34, Sec. 1.8]. A dtmc is said to be ergodic if it is irreducible, positive recurrent, and also aperiodic, i.e., the chain does not become stuck in a cycle of fixed length. For an ergodic dtmc, the long-term expected value and the expected value of the time average converge to the stationary mean [34, Sec. 1.10]:
@@
Unless stated otherwise, the Markov chains induced by the update policies in this section are ergodic. Since every VAoI state can reach state 0 (or 111) and vice versa with positive probability, and the transitions allow for self-resets and exits from any loop, the chain is irreducible, positive recurrent, and aperiodic. We proceed with analyzing their steady-state distributions.
@@ IV-B VAoI of uniform policy
Here, $\mu_{n}^{(q)}$ denotes the steady-state probability of the time-homogeneous Markov chain: $Y^{(q)}(k)=\Delta(kD+q)$, which samples the original VAoI process at time indices spaced at intervals of $D$, starting from phase offset $q\in\{1,2,\dots,D\}$. This $\mu_{n}$ represents the long-run time average (or steady-state occupancy probability) for the original periodically time-inhomogeneous Markov process [34, Sec. 1.8, Theorem 1.8.5].
@@ IV-C VAoI of threshold policy
TABLE I: Average VAoI at node $N\!+\!1$ for various update policies.
@@ V Analysis of VAoI in a Multi-hop Setup
The VAoI at node $i\!+\!1$ is given by:
@@
The variable $\tau_{N}$ is the relaying delay of each version from node 111 to node $N\!+\!1$ through $N$ relaying nodes, while $\beta_{N}$ represents the number of version generations at the source during this delay. Specifically, $\tau_{N}$ and $\beta_{N}$ are the sums of independent Geometric and Binomial RVs, $\{m_{i}\}_{i=1}^{N}$ and $\{\eta_{m_{i}}\!\mid\!m_{i}\}_{i=1}^{N}$, respectively. Their probability mass functions (PMFs) can be derived by convolving the PMFs of the individual components. This derivation simplifies in two cases: (1) when $\{\rho_{i}\}_{i=1}^{N}\!=\!\rho$, $\tau_{N}$ follows a Negative Binomial distribution, $\tau_{N}\!\sim\!NegBin(N,\rho)$, representing the number of trials required to achieve the $N$-th success, with $\mathbb{P}(\tau_{N}\!=\!\ell)\!=\!\binom{\ell\!-\!1}{N\!-\!1}\rho^{N}(1\!-\!\rho)^{\ell\!-\!N}$ for $\ell\!=\!N,N\!+\!1,\dots$; and (2) when $N$ is large, by the Central Limit Theorem, $\tau_{N}\!\sim\!\mathcal{N}\!\left(\sum_{i=1}^{N}\!\!\frac{1}{\rho_{i}},\,\sum_{i=1}^{N}\!\!\frac{1-\rho_{i}}{\rho_{i}^{2}}\right)$. Moreover, since $\{m_{i}\}_{i=1}^{N}$ are independent, $\eta_{\tau_{N}}\!\mid\!\tau_{N}\sim Bin(\tau_{N},p_{g})$ always holds.
@@ VI Numerical Results
We evaluate and discuss the analytical results, starting with the single-hop setup under various update policies and extending the analysis to the multi-hop setup. Simulations were conducted over $10^{4}$ time slots, and the results were averaged over 400400400 Monte Carlo iterations to obtain the steady-state values.
@@ VI-A Single-hop Setup
Figure 6: Stationary distribution of VAoI at node 111 under various policies for $\alpha=0.25$.
Figure 7: Stationary distribution of VAoI at node 111 under various policies for $\alpha=0.05$.
Figure 8: Average VAoI at node 111 under various policies for $\alpha=0.05$.
[…]
@@
Figure 11: Average VAoI at node $N\!+\!1$ under various policies ($\alpha=0.05$, $N=6$ and 242424).
@@ VI-B Multi-hop Setup
According to Lemma 3, the VAoI at the destination node $N+1$ in a multi-hop setup with $N$ relays equals the average VAoI at node 111, shifted by the relaying delay $\tau_{N}$, plus the number of version generations at the source during this delay, denoted by the Binomial RV $\beta_{N}$. Consequently, the long-term average VAoI, $\bar{\Delta}_{N+1}$, equals the average VAoI at node 111, $\bar{\Delta}_{1}$, plus an offset corresponding to the expected number of version generations during relaying, given by $p_{g}\sum_{i=1}^{N}\frac{1}{\rho_{i}}$. The relaying delay $\tau_{N}$ follows a negative Binomial distribution under ${\rho_{i}}_{i=1}^{N}=\rho$. As $N$ increases, this distribution converges to a normal distribution, as illustrated in Fig. 11 for $\rho=0.7$ with $N=6$ and 242424. The average VAoI, $\bar{\Delta}_{N+1}$, is depicted in Fig. 11 for various policies as a function of $\rho$, with parameters $p_{s}=0.8$, $p_{g}=0.3$, $\alpha=0.05$, for $N=6$ and 242424. The average VAoI, $\bar{\Delta}_{N+1}$, exhibits a linear increase with respect to $N$ and a reciprocal polynomial decrease with respect to $\rho$, such that $\bar{\Delta}\propto\frac{N}{\rho}$. That is, as the network size increases, maintaining the average VAoI at the destination requires the incorporation of more reliable links.

### v2

@@ Abstract
Timely and informative data dissemination in communication networks is essential for enhancing system performance and energy efficiency, as it reduces the transmission of outdated or redundant data. Timeliness metrics, such as Age of Information (AoI), effectively quantify data freshness; however, these metrics fail to account for the intrinsic informativeness of the content itself. To address this limitation, content-based metrics have been proposed that combine both timeliness and informativeness. Nevertheless, existing studies have predominantly focused on evaluating average metric values, leaving the complete distribution—particularly in multi-hop network scenarios—largely unexplored. In this paper, we provide a comprehensive analysis of the stationary distribution of the Version Age of Information (VAoI), a content-based metric, under various scheduling policies, including randomized stationary, uniform, and threshold-based policies, with transmission constraints in single-hop and multi-hop networks. We derive closed-form expressions for the stationary distribution and average VAoI under these scheduling approaches. Furthermore, for threshold-based scheduling, we analytically determine the optimal threshold value that minimizes VAoI and derive the corresponding optimal VAoI in closed form. Numerical evaluations verify our analytical findings and provide valuable insights into leveraging VAoI in the design of efficient communication networks.
@@ I Introduction
Efficient data management is a critical requirement for ensuring optimal performance in communication networks across a wide range of applications, from single-hop IoT monitoring systems to multi-hop satellite-based networks. As the volume of data generated by these networks increases, transmitting all data indiscriminately, without considering its semantic significance or task-specific utility, becomes increasingly unsustainable. Such an approach results in excessive consumption of critical resources, including energy and bandwidth, ultimately compromising system practicality and degrading overall performance. To address these challenges, there is an urgent need for network management approaches that optimize data transmission by leveraging goal-oriented semantic metrics; that is, by delivering the most timely and informative data within a constrained frequency of data transmissions [27, 28].
The AoI [25] is a widely used semantic metric that quantifies the freshness of information in status update systems as the time elapsed since the generation of the most recently received data. AoI-aware scheduling effectively minimizes staleness by adapting transmissions to source data arrivals and network service times [39]. However, AoI captures freshness solely through data timestamps, without accounting for actual changes in the source content. As a result, simply refreshing timestamps may fail to deliver new information, and avoiding such redundant updates can reduce data transmission and energy consumption.
To address this limitation, content-based metrics such as Age of Incorrect Information (AoII) [29] and VAoI [41] have been introduced. AoII adds a distortion-aware dimension by measuring the staleness of incorrect information—specifically when the receiver’s content deviates from the source—unlike AoI, which treats both correct and incorrect data uniformly. However, AoII requires precise knowledge of the content or state of information at both source and destination for comparison, which is practical only when the state space is small and fully modeled, with all transitions known. In many real-world applications, such complete knowledge of source content and transitions may not be available. In such cases, VAoI offers a more practical, content-based metric by focusing solely on content changes at the source, where data evolve through successive, non-reverting versions. This requires minimal knowledge: at any time, either a new or the previous version exists, and the receiver must track these versions as timely as possible. Defined as the number of versions by which the receiver lags behind the source, VAoI further improves upon AoI by replacing timestamps with version numbers, thereby eliminating the challenging requirement of clock synchronization between the transmitter and receiver [33]. It is computed simply by comparing the receiver’s stored version with the source’s current version.
@@ II Related Works
Several studies have investigated the distributions of AoI and Peak AoI (PAoI) in continuous-time systems using queueing theory [15, 11, 20, 40, 13, 22, 1, 30, 18, 3, 19], while others have examined discrete-time settings [26, 42, 2, 21, 43]. Notably, [26] derives general expressions for the stationary distributions and generating functions of AoI and PAoI in discrete-time single-server queues under various disciplines, together with methods for evaluating nonlinear age functions. Extending stochastic hybrid system techniques to discrete time, [42] models AoI and packet age as a two-dimensional Markov process in bufferless queues with Bernoulli arrivals. A matrix-analytic framework based on quasi-birth–death processes is proposed in [2] to obtain exact per-source AoI and PAoI distributions in multi-source IoT systems with discrete phase-type service times. The study in [21] investigates age-optimal packet scheduling with delayed feedback and long-term resource constraints, providing closed-form benchmarks for random and deterministic policies; and [43] analyzes AoI and PAoI in multi-source Ber/Geo/1/1 systems under preemptive and non-preemptive policies, deriving closed-form expressions for distributions and averages.
@@
Regarding multi-hop networks, [6] derives the distribution of discrete-time AoI in $N$-hop systems with time-invariant packet loss through recursive formulations, while [37] studies continuous-time AoI and PAoI distributions in two-hop scenarios. However, most existing research on multi-hop networks, including [35, 7, 10, 14, 36, 24, 34, 5, 17], primarily focuses on average metrics.
@@ III System Model
We consider a communication network for transmitting data from a source to a destination node, which may be separated by either a single hop or multiple hops. The data at the source are constantly sampled from an information source and then transmitted according to an update policy. The update policy schedules each data sample or update to be either transmitted or skipped, while satisfying a constraint on the long-term average transmission rate. Specifically, the average update rate must not exceed a predefined limit. We assume a slotted time axis $t\in\{0,1,2,\cdots\}$, and our objective is to investigate the distribution of the VAoI at the destination node in both single-hop and multi-hop scenarios under various rate-constrained policies. The details of the system model are explained below.
@@
Version Age of Information: We adopt the VAoI as the performance metric that captures both the timeliness and relevance of information. Unlike the AoI, which measures the time elapsed since the generation timestamp $u(t)$ of the freshest received update, $\Delta^{AoI}(t)\!=\!t\!-\!u(t),$ the VAoI quantifies how many versions the receiver lags behind the information source. At time slot $t$, the VAoI, denoted $\Delta(t)$, is defined as $\Delta(t)\!=\!V_{S}(t)\!-\!V_{R}(t)$, where $V_{S}(t)$ is the version index at the information source and $V_{R}(t)$ the version stored at the receiver. Fig. 4 shows the evolution of AoI and VAoI over discrete slots. A successful update occurs at t=3t\!=\!3, causing AoI and VAoI to drop to $1$ and 0, respectively, since the update is one slot old and no new versions have been generated. Between t=3t\!=\!3 and the next update at t=12t\!=\!12, the AoI grows linearly to $9$, while the VAoI reaches $4$, reflecting the generation of $4$ new versions at the information source.
Fig. 5 illustrates VAoI dynamics in a two-hop network. The information source generates new versions at time slots 0, 22, 33, 66, and $8$, with its version index $V_{S}(t)$ increasing by one at the start of each subsequent slot. Node 0 holds the most recent version available in the network; hence, its VAoI is always zero, i.e., $V_{0}(t)\!=\!V_{S}(t)$. The versions at node $1$, denoted by $V_{1}(t)$, are updated according to the update policy and the transmissions from node 0. If a transmission at time $t$ succeeds, then $V_{1}(t\!+\!1)\!=\!V_{0}(t)$; otherwise, $V_{1}(t\!+\!1)\!=\!V_{1}(t)$. The corresponding VAoI, $\Delta_{1}(t)\!=\!V_{S}(t)\!-\!V_{1}(t)$, is listed in parentheses in the second row of the table and plotted in blue. Similarly, node $2$ stores versions $V_{2}(t)$ received from node $1$, which transmits in every slot. [^1] In practice, if feedback channels are available between relay nodes, retransmissions of already delivered versions can be avoided; otherwise, relay nodes transmit continuously each time slot. In both cases, however, the presented VAoI analysis remains valid. Here, $V_{2}(t\!+\!1)\!=\!V_{1}(t)$ if the transmission at $t$ succeeds; otherwise, $V_{2}(t\!+\!1)\!=\!V_{2}(t)$. The VAoI at node $2$, $\Delta_{2}(t)\!=\!V_{S}(t)\!-\!V_{2}(t)$, is shown in parentheses in the third row of the table and plotted in green in Fig. 5.
@@ IV Analysis of VAoI in a Single-hop Setup
We analyze the VAoI in a single-hop system under the three aforementioned update policies using a dtmc model. The VAoI in the next time slot is stochastically determined by the current VAoI, the update policy, and system dynamics, including the version generation and channel success processes.
@@
where $\sum_{n=0}^{\infty}\mu_{n}=1$. Here, $P_{jn}$ denotes the transition probability from state $j$ to state $n$, i.e., $P_{jn}=\mathbb{P}\left(\Delta(t+1)=n\mid\Delta(t)=j\right)$. The steady-state distribution of a dtmc exists if it is irreducible, i.e., every state can be reached from every other state, and positive recurrent, i.e., the expected return time to each state is finite [31, Sec. 1.8]. A dtmc is said to be ergodic if it is irreducible, positive recurrent, and also aperiodic, i.e., the chain does not become stuck in a cycle of fixed length. For an ergodic dtmc, the long-term expected value and the expected value of the time average converge to the stationary mean [31, Sec. 1.10]:
@@
Unless stated otherwise, the Markov chains induced by the update policies in this section are ergodic. Since every VAoI state can reach state 0 (or $1$) and vice versa with positive probability, and the transitions allow for self-resets and exits from any loop, the chain is irreducible, positive recurrent, and aperiodic. We proceed with analyzing their steady-state distributions.
@@ IV-B VAoI of uniform policy
Here, $\mu_{n}^{(q)}$ denotes the steady-state probability of the time-homogeneous Markov chain: $Y^{(q)}(k)=\Delta(kD+q)$, which samples the original VAoI process at time indices spaced at intervals of $D$, starting from phase offset $q\in\{1,2,\dots,D\}$. This $\mu_{n}$ represents the long-run time average (or steady-state occupancy probability) for the original periodically time-inhomogeneous Markov process [31, Sec. 1.8, Theorem 1.8.5].
@@ IV-C VAoI of threshold policy
TABLE I: Average VAoI at node N+1N\!+\!1 for various update policies.
@@ V Analysis of VAoI in a Multi-hop Setup
The VAoI at node i+1i\!+\!1 is given by:
@@
The variable $\tau_{N}$ is the relaying delay of each version from node $1$ to node N+1N\!+\!1 through $N$ relaying nodes, while $\beta_{N}$ represents the number of version generations at the source during this delay. Specifically, $\tau_{N}$ and $\beta_{N}$ are the sums of independent Geometric and Binomial RVs, $\{m_{i}\}_{i=1}^{N}$ and $\{\eta_{m_{i}}\!\mid\!m_{i}\}_{i=1}^{N}$, respectively. Their probability mass functions (PMFs) can be derived by convolving the PMFs of the individual components. This derivation simplifies in two cases: (1) when $\{\rho_{i}\}_{i=1}^{N}\!=\!\rho$, $\tau_{N}$ follows a Negative Binomial distribution, $\tau_{N}\!\sim\!NegBin(N,\rho)$, representing the number of trials required to achieve the $N$-th success, with $\mathbb{P}(\tau_{N}\!=\!\ell)\!=\!\binom{\ell\!-\!1}{N\!-\!1}\rho^{N}(1\!-\!\rho)^{\ell\!-\!N}$ for $\ell\!=\!N,N\!+\!1,\dots$; and (2) when $N$ is large, by the Central Limit Theorem, $\tau_{N}\!\sim\!\mathcal{N}\!\left(\sum_{i=1}^{N}\!\!\frac{1}{\rho_{i}},\,\sum_{i=1}^{N}\!\!\frac{1-\rho_{i}}{\rho_{i}^{2}}\right)$. Moreover, since $\{m_{i}\}_{i=1}^{N}$ are independent, $\beta_{N}\!\mid\!\tau_{N}\sim Bin(\tau_{N},p_{g})$ always holds.
@@ VI Numerical Results
We evaluate the analytical results, starting with the single-hop setup under various update policies and extending the analysis to the multi-hop setup. Simulations were conducted over $10^{4}$ time slots, and the results were averaged over $400$ Monte Carlo iterations to obtain the steady-state values.
@@ VI-A Single-hop Setup
Figure 6: Stationary distribution of VAoI at node 11 under various policies for $\alpha=0.25$.
Figure 7: Stationary distribution of VAoI at node 11 under various policies for $\alpha=0.05$.
Figure 8: Average VAoI at node 11 under various policies for $\alpha=0.05$.
Fig. 8 illustrates that, to maintain a target VAoI of, for instance, $2$, the uniform policy reduces the required update rate by $35\%$ (from $0.188$ to $0.121$), while the optimal threshold policy achieves a $54\%$ reduction (to $0.086$), compared to the randomized policy. These results highlight that optimal VAoI-aware policies can significantly reduce the transmission rate without compromising the conveyed information, thereby improving energy efficiency in communication networks.
@@
Figure 11: Average VAoI at node N+1N\!+\!1 under various policies ($\alpha=0.05$, $N=6$ and $24$).
@@ VI-B Multi-hop Setup
According to Lemma 3, the VAoI at the destination node $N+1$ in a multi-hop setup with $N$ relays equals the average VAoI at node $1$, shifted by the relaying delay $\tau_{N}$, plus the number of version generations at the source during this delay, denoted by the Binomial RV $\beta_{N}$. Consequently, the long-term average VAoI, $\bar{\Delta}_{N+1}$, equals the average VAoI at node $1$, $\bar{\Delta}_{1}$, plus an offset corresponding to the expected number of version generations during relaying, given by $p_{g}\sum_{i=1}^{N}\frac{1}{\rho_{i}}$. The relaying delay $\tau_{N}$ follows a negative Binomial distribution under ${\rho_{i}}_{i=1}^{N}=\rho$. As $N$ increases, this distribution converges to a normal distribution, as illustrated in Fig. 11 for $\rho=0.7$ with $N=6$ and $24$. The average VAoI, $\bar{\Delta}_{N+1}$, is depicted in Fig. 11 for various policies as a function of $\rho$, with parameters $p_{s}=0.8$, $p_{g}=0.3$, $\alpha=0.05$, for $N=6$ and $24$. The average VAoI, $\bar{\Delta}_{N+1}$, exhibits a linear increase with respect to $N$ and a reciprocal polynomial decrease with respect to $\rho$, such that $\bar{\Delta}\propto\frac{N}{\rho}$. That is, as the network size increases, maintaining the average VAoI at the destination requires the incorporation of more reliable links.